1. Ashley sells a hammer for $ 50. Its variable expense is $ 35 per unit and its fixed expense is $ 60,000. The current volume is 5,000 units. Each question is independent.

a. Find the CM ratio, the degree of operating leverage, and the break-even point in dollars.

b. Compute the percentage change of operating income when the sales for the next period increase by 25%.

c. If the variable expense decreases by $ 5, find the CM ratio and the break-even sales in the unit.

d. When the fixed expense changes to $ 80,000, find the sales volume to maintain the current operating income.

e. When the fixed expense decreases to $ 40,000 and the variable expense increases to $ 42, how many units should be sold to attain $ 80,000 of operating income?
Answer #1

a) CM ratio = 15/50 = 30%

Operating leverage = Contribution margin/Operating income = 75000/15000 = 5 Times

Break even sales = 60000/.30 = $200000

b) Operating income increase by = 25*5 = 125%

c) CM ratio = 20/50 = 40%

Break even unit = 60000/20 = 3000 Units

d) Sales volume = (80000+15000)/15 = 6333 Units

e) Required unit = (40000+80000)/8 = 15000 Units